Suppose market demand andmarket supply are given by Qd = 15 –4P and Qs = -3+2P What are the equilibrium quantity and price in this market? Show your work!!!
Answer #1

First find the equilibrium price

Qd = Qs

15 - 4P = -3 + 2P

15 + 3 = 2P + 4P

18 = 6P

P* = $3 per unit

Qs = -3 + 2*3 = 3 units and Qd = 15 - 4*3 = 3 units

Hence the equilibrium quantity is 3 units and the equilibrium price is $3 per unit in this market